# A quasilinear elliptic equation with absorption term and Hardy potential

### v1

@@ 5.5 Harnack inequality
From Trudinger [26, p.724], there exists a constant $C_{1}$ depending on $N,p,\left|x_{0}\right|\sup\left|\varphi\right|_{B(x_{0},\frac{\left|x_{0}\right|}{2})}$ such that
@@
From (5.1), $\left|x_{0}\right|\sup\left|\psi\right|_{B(x_{0},\frac{\left|x_{0}\right|}{2})}$ is bounded by a constant only depending on $N,p,q,\theta;$ then $C$ only depends on $N,p,q,\theta,\mu.$ Then (5.5) holds by connecting two points $x_{1},x_{2}$ such that $\left|x_{1}\right|=\left|x_{2}\right|=r<\frac{r_{0}}{2}$ by 10 connected balls of radius $\frac{r}{6}.$
@@ 5.7 Precise convergence results
$u_{r}(\xi)=\frac{u(r\xi)}{v_{i}(r)}\qquad\text{for }0<\left|\xi\right|<\frac{r_{0}}{r}.$
@@
$-\Delta_{p}w+\mu\frac{w^{p-1}}{\left|\xi\right|^{p}}=0\text{ in }\mathbb{R}^{N}\backslash\left\{0\right\};$
@@
We get analogous results near $\infty$,∞ , by using Proposition 5.6 in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ with $\delta=S_{i}>\gamma$.
@@
$\lim_{x\longrightarrow 0}\left|x\right|^{-\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{p}}=k\text{ (resp. }\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{-\frac{N-p}{p}}(\ln\left|x\right|)^{\frac{2}{p}}=k\text{)}$
@@
such that $\lim_{r\longrightarrow 0}r^{\frac{N-p}{p}}\left|\ln r\right|^{-1}v_{0}(r)=1,$ given at Lemma 4.5. Let
@@
Since $\frac{N-p}{p}<\gamma$ there holds $\lim_{r\longrightarrow 0}(r^{\gamma}v_{0})^{q+1-p}=0$. By hypothesis, for $u(x)\leq Cv_{0}\left(\left|x\right|\right)\leq C_{1}\left|x\right|^{-\frac{N-p}{p}}\left|\ln\left|x\right|\right|$ in $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\}.$ From Remark 5.7, for any $x,x^{\prime}\in$ $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\},$
$\left|\nabla u(x)\right|\leq C_{2}\left|x\right|^{-\frac{N}{p}}\left|\ln\left|x\right|\right|$
$\left|\nabla u(x)-\nabla u(x^{\prime})\right|\leq C_{3}(\left|x^{\prime}\right|^{-\frac{N-p}{p}}+\left|x\right|^{-\frac{N-p}{p}})\left|x\right|^{-(1+\alpha)}\left|\ln\left|x\right|\right|\left|x-x^{\prime}\right|^{\alpha}.$
then for $\left|\xi\right|,\left|\xi^{\prime}\right|<\frac{r_{0}}{2r}<\frac{1}{2r}$
$u_{r}(\xi)\leq C\frac{v_{0}(r\xi)}{v_{0}(r)}\leq C_{1}\left|\xi\right|^{-\frac{N-p}{p}}\frac{\left|\ln r\xi\right|}{\left|\ln r\right|}\leq C_{1}\left|\xi\right|^{-\frac{N-p}{p}}(1+\frac{\left|\ln\xi\right|}{\left|\ln r\right|})$
$\left|\nabla u_{r}(\xi)\right|=\frac{r}{v_{0}(r)}\left|\nabla u(r\xi)\right|\leq C_{4}r\left|r\xi\right|^{-\frac{N}{p}}\frac{\left|\ln r\xi\right|}{r^{-\frac{N-p}{p}}\left|\ln r\right|}=C_{4}\left|\xi\right|^{-\frac{N}{p}}\frac{\left|\ln r\xi\right|}{\left|\ln r\right|}$
$\left|\nabla u_{r}(\xi)-\nabla u_{r}(\xi^{\prime})\right|\leq C_{3}r^{\alpha}(\left|\xi^{\prime}\right|^{-\frac{N-p}{p}}+\left|\xi\right|^{-\frac{N-p}{p}})\left|\xi\right|^{-(1+\alpha)}\left|\ln r\xi\right|\left|\xi-\xi^{\prime}\right|^{\alpha}.$
@@
in $\mathbb{R}^{N}\backslash\left\{0\right\},$ and $\xi_{r_{n}}\longrightarrow\xi_{0}\in S^{n-1},$ $u_{r_{n}}(\xi_{r_{n}})\longrightarrow w(\xi_{0})$ and $\widetilde{k}(r_{n})=\frac{u_{r_{n}}(\xi_{r_{n}})}{v_{0}(r_{n})}$. But from [14, Example 1.1], the unique solutions of this equation are the functions $\xi\longmapsto\ell\left|\xi\right|^{-\frac{N-p}{p}},$ $\ell>0.$ Then $w(\xi_{0})=\ell\left|\xi_{0}\right|^{-\frac{N-p}{p}}=\ell$ and $\widetilde{k}(r_{n})=u_{r_{n}}(\xi_{r_{n}})\longrightarrow k.$ Then $\ell=k,$ thus $w$ is independant of the choice of $r_{n},$ then $\lim_{r\longrightarrow 0}u(r\xi)=k$ any $\xi\in\mathbb{R}^{N}\backslash\left[0\right]$, in particular for $\xi\in S^{n-1}.$ Then $\lim_{x\longrightarrow 0}\frac{u}{v_{0}}(x)=k.$ We get similar results in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ when $\frac{N-p}{p}>\gamma$
@@ 6 Proof of the nonradial results
Proof of Theorem 2.1. (i) Let $u$ be a solution in $B_{r_{0}}\backslash\left\{0\right\}.$ From Theorem 5.10, there exists radial functions such that $v\leq u\leq w\leq cv$ in $B_{r_{0}}\backslash\left\{0\right\}.$ First suppose $\mu>\mu_{0}.$ From Theorem 4.6, either $\lim_{r\longrightarrow 0}r^{\gamma}v=a^{\ast},$ then also $\lim_{r\longrightarrow 0}r^{\gamma}w=a^{\ast},$ then by squeezing $\lim_{x\longrightarrow 0}\left|x\right|^{\gamma}u=a^{\ast}$; or $\lim_{r\longrightarrow 0}r^{S_{i}}v=k_{i}^{\prime}>0,$ with $i=1$ or $2$, then also $\lim_{r\longrightarrow 0}r^{S_{i}}w=k_{i}^{\prime\prime}>0,$ implying that $u\asymp$ $\left|x\right|^{-S_{i}},$ then $\lim_{x\longrightarrow 0}\left|x\right|^{S_{i}}u=k_{i}>0$ from Proposition 5.11. For $\mu=\mu_{0}$ we still have a possibility that $\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}(\left|\ln\right|^{-\frac{2}{p}}\left|x\right|)v(\left|x\right|)=\ell_{1}>0$, and $\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}(\left|\ln\right|^{-\frac{2}{p}}\left|x\right|)w(\left|x\right|)=\ell_{2}>0$, then we get (2.17) from Proposition 5.12.
(ii) Let $u$ be a solution in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}.$ Then from 4.6, $\lim_{r\longrightarrow\infty}r^{\gamma}v=a^{\ast}=\lim_{r\longrightarrow\infty}r^{\gamma}w,$ thus $\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{\gamma}u=a^{\ast}$.
(iii) Let $u$ be a solution in $\mathbb{R}^{N}\backslash\left\{0\right\}$. Then $v\leq u\leq w\leq cv$ in $\mathbb{R}^{N}\backslash\left\{0\right\}$. From Theorem 4.6, all the behaviours of radial solutions in $\mathbb{R}^{N}\backslash\left\{0\right\}$ are still described. Either $\mu>\mu_{0}$ and $r^{\gamma}v\equiv r^{\gamma}v=a^{\ast},$ thus $\left|x\right|^{\gamma}u\equiv a^{\ast}$. Or $\mu>\mu_{0}$, and $\lim_{r\longrightarrow 0}r^{S_{1}}v=k_{i}^{\prime}>0$, and $\lim_{r\longrightarrow\infty}r^{\gamma}v=a^{\ast}$; then $\lim_{x\longrightarrow 0}\left|x\right|^{S_{1}}u=k_{i}>0$ from Proposition 5.11, and $\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{\gamma}u=a^{\ast}$. For given $k_{1}>0,$ there exists a unique function $u$ satisfying these two conditions. Indeed if $\widetilde{u}$ is another such solution, then $(1+\varepsilon)\widetilde{u}$ is a supersolution, and near 00 and $\infty$,∞ , it is greater than $u,$ then $(1+\varepsilon)\widetilde{u}\geq u;$ then $\widetilde{u}\geq u,$ and $\widetilde{u}=u.$ By uniqueness $u$ is radial. Finally if $\mu=\mu_{0},$ then $\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}(\left|\ln\right|^{-\frac{2}{p}}\left|x\right|)u(x)=\ell>0$ from Proposition 5.12, and $\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{\gamma}u(x)=a^{\ast}.$ By comparison as above, it is unique and radial.
@@ 7.1.1 Behaviour near the fixed point $\mathbf{M}_{0}$
are the roots of the trinom $T(\lambda);$ and the product of the roots is negative, thus $\mathbf{M}_{0}$ is saddle point. The eigenvectors can be computed easily.
@@ 7.1.2 Behaviour near the fixed points $\mathbf{A}_{1}$,$\mathbf{A}_{2}$
$\det(\begin{array}[]{cc}pS_{i}-N+p-\lambda&-1\\ 0&(q+1-p)(\gamma-S_{i})-\lambda\end{array})=0,$
@@ 7.2 Proofs and comments
then $\mathcal{T}_{1}$ lies in $\mathcal{R}_{1}$ as $t\rightarrow\infty$, then it stays in it, and necessarily converges to $\mathbf{A}_{1}$ (even in the case $\rho_{1}=\eta_{1})$ So there is a, unique trajectory, namely $\mathcal{T}_{1}$, joining $\mathbf{A}_{1}$ to $\mathbf{M}_{0}$, then $u$ satisfies (4.13).
$\bullet$ The trajectory $\mathcal{T}_{3}$ starts from $\mathbf{M}_{0}$ in the region $\mathcal{R}_{3};$ and $\mathcal{R}_{3}$ is positively invariant: indeed the vector field is entering $\mathcal{R}_{3}$ at any point or the curve $\mathcal{C}_{2}$, except possibly at the singular point $(0,\left|\mu\right|)$ when $\mu<0$ and $p>2$. But no trajectory with a point in $\mathcal{R}_{3}$ can pass by this point, since such trajectories satisfy $G<0$, from then $\mathcal{T}_{3}$ stays in $\overline{\mathcal{R}_{3}}$.. Consider any corresponding solution $u(r)=U(t)$ defined on an maximal interval $(-\infty,T)$. Suppose that $V$ is bounded; then $F(G)$ is bounded, so $\left|G\right|$ is bounded, then $(G,V)$ converges to some $(\ell,\Lambda)$, then $(G_{t},V_{t})$ has a finite limit. It cannot happen that $(\ell,\Lambda)=(0,-\mu)$, because we have seen that the trajectories passing by this point are contained in the set $G<0$. Then $T=\infty$, thus $(G,V)$ converges to a fixed point; it is impossible, since there is no other fixed point in $\overline{\mathcal{R}_{3}}$. Then $T$ is finite, and $\lim_{t\longrightarrow T}V=\infty;$ if $\lambda=\lim_{t\longrightarrow T}G$ is finite, and $\lim_{t\longrightarrow T}\frac{V_{t}}{V}=c>0$, which contradicts $\lim_{t=T}V=\infty$. Then $\lim_{t\longrightarrow T}G=-\infty$. It implies that the function $u$, such that $u\sim_{r\longrightarrow 0}u^{\ast}$decreasing as $r\rightarrow 0$, has a minimum point, and then is increasing to $\infty$ as $r\rightarrow e^{T}$. Namely we obtain a solution $u$, satisfying (4.17), often called large solution.
@@
for some $\tau\in\left(0,1\right),$ thus $F_{1}(G_{1}+\overline{G})\sim_{t\longrightarrow-\infty}c\overline{G}^{2}$ with $c=\frac{1}{2}\frac{p}{p-1}S_{1}^{2-p};$ and
$\overline{G}_{t}=c\overline{G}^{2}(1+O(\overline{G}))-V$
@@
$\bullet$ In the general case $p>1$, we can solve the equation (7.13) by quadrature, and get

### v2

@@ 5.5 Harnack inequality
From Trudinger [26, p.724], there exists a constant $C_{1}$ depending on $N,p,\left|x_{0}\right|\sup_{B(x_{0},\frac{\left|x_{0}\right|}{2})}\left|\varphi\right|$ such that
@@
From (5.1), $\left|x_{0}\right|\sup_{B(x_{0},\frac{\left|x_{0}\right|}{2})}\left|\psi\right|$ is bounded by a constant only depending on $N,p,q,\theta;$ then $C$ only depends on $N,p,q,\theta,\mu.$ Then (5.5) holds by connecting two points $x_{1},x_{2}$ such that $\left|x_{1}\right|=\left|x_{2}\right|=r<\frac{r_{0}}{2}$ by 10 connected balls of radius $\frac{r}{6}.$
@@ 5.7 Precise convergence results
$u_{r}(\xi)=\frac{u(r\xi)}{v_{i}(r)}\qquad\text{for }0<\left|\xi\right|<\frac{r_{0}}{r}$
@@
$-\Delta_{p}w+\mu\frac{w^{p-1}}{\left|\xi\right|^{p}}=0\text{ in }\mathbb{R}^{N}\backslash\left\{0\right\},$
@@
We get analogous results near $\infty$,∞ , by using Proposition 5.6 in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ with $\delta=S_{i}>\gamma$.$\varpi$
@@
$\lim_{x\longrightarrow 0}\left|x\right|^{-\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{-\frac{2}{p}}u=k\text{ (resp. }\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{-\frac{N-p}{p}}(\ln\left|x\right|)^{-\frac{2}{p}}u=k\text{)}$
@@
such that $\lim_{r\longrightarrow 0}r^{\frac{N-p}{p}}\left|\ln r\right|^{-\frac{2}{p}}v_{0}(r)=1,$ given at Lemma 4.5. Let
@@
Since $\frac{N-p}{p}<\gamma$ there holds $\lim_{r\longrightarrow 0}(r^{\gamma}v_{0})^{q+1-p}=0$. By hypothesis, for $u(x)\leq Cv_{0}\left(\left|x\right|\right)\leq C_{1}\left|x\right|^{-\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{p}}$ in $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\}.$ From Proposition 5.7, for any $x,x^{\prime}\in$ $B_{\frac{r_{0}}{2}}\backslash\left\{0\right\},$
$\left|\nabla u(x)\right|\leq C_{2}\left|x\right|^{-\frac{N}{p}}(\left|\ln\left|x\right|\right|+1)^{\frac{2}{p}}\leq 2^{\frac{2}{p}}C_{2}\left|x\right|^{-\frac{N}{p}}\left|\ln\left|x\right|\right|^{\frac{2}{p}}$
$\left|\nabla u(x)-\nabla u(x^{\prime})\right|\leq 2^{\frac{2}{p}}C_{3}(\left|x^{\prime}\right|^{-\frac{N-p}{p}}+\left|x\right|^{-\frac{N-p}{p}})\left|x\right|^{-(1+\alpha)}\left|\ln\left|x\right|\right|^{\frac{2}{p}}\left|x-x^{\prime}\right|^{\alpha}.$
Then for $\left|\xi\right|,\left|\xi^{\prime}\right|<\frac{r_{0}}{2r}<\frac{1}{2r},$
$u_{r}(\xi)\leq C\frac{v_{0}(r\xi)}{v_{0}(r)}\leq C_{1}\left|\xi\right|^{-\frac{N-p}{p}}(\frac{\left|\ln r\xi\right|}{\left|\ln r\right|})^{\frac{2}{p}}\leq C_{1}\left|\xi\right|^{-\frac{N-p}{p}}(1+\frac{\left|\ln\xi\right|}{\left|\ln r\right|})^{\frac{2}{p}}$
$\left|\nabla u_{r}(\xi)\right|=\frac{r}{v_{0}(r)}\left|\nabla u(r\xi)\right|\leq C_{4}r\left|r\xi\right|^{-\frac{N}{p}}\frac{\left|\ln r\xi\right|^{\frac{2}{p}}}{r^{-\frac{N-p}{p}}\left|\ln r\right|^{\frac{2}{p}}}=C_{4}\left|\xi\right|^{-\frac{N}{p}}(1+\frac{\left|\ln\xi\right|}{\left|\ln r\right|})^{\frac{2}{p}}$
$\left|\nabla u_{r}(\xi)-\nabla u_{r}(\xi^{\prime})\right|\leq C_{3}r^{\alpha}(\left|\xi^{\prime}\right|^{-\frac{N-p}{p}}+\left|\xi\right|^{-\frac{N-p}{p}})\left|\xi\right|^{-(1+\alpha)}(1+\frac{\left|\ln\xi\right|}{\left|\ln r\right|})^{\frac{2}{p}}\left|\xi-\xi^{\prime}\right|^{\alpha}.$
@@
in $\mathbb{R}^{N}\backslash\left\{0\right\},$ and $\xi_{r_{n}}\longrightarrow\xi_{0}\in S^{n-1},$ $u_{r_{n}}(\xi_{r_{n}})\longrightarrow w(\xi_{0})$ and $\widetilde{k}(r_{n})=\frac{u_{r_{n}}(\xi_{r_{n}})}{v_{0}(r_{n})}$. But from [14, Example 1.1], the only solutions of this equation are the functions $\xi\longmapsto\ell\left|\xi\right|^{-\frac{N-p}{p}},$ $\ell>0.$ Then $w(\xi_{0})=\ell\left|\xi_{0}\right|^{-\frac{N-p}{p}}=\ell$ and $\widetilde{k}(r_{n})=u_{r_{n}}(\xi_{r_{n}})\longrightarrow k.$ Then $\ell=k,$ thus $w$ is independant of the choice of $r_{n},$ then $\lim_{r\longrightarrow 0}u(r\xi)=k$ any $\xi\in\mathbb{R}^{N}\backslash\left[0\right]$, in particular for $\xi\in S^{n-1}.$ Then $\lim_{x\longrightarrow 0}\frac{u}{v_{0}}(x)=k.$ We get similar results in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}$ when $\frac{N-p}{p}>\gamma$
@@ 6 Proof of the nonradial results
Proof of Theorem 2.1. (i) Let $u$ be a solution in $B_{r_{0}}\backslash\left\{0\right\}.$ From Theorem 5.10, there exists radial functions such that $v\leq u\leq w\leq cv$ in $B_{r_{0}}\backslash\left\{0\right\}.$ First suppose $\mu>\mu_{0}.$ From Theorem 4.6, either $\lim_{r\longrightarrow 0}r^{\gamma}v=a^{\ast},$ then also $\lim_{r\longrightarrow 0}r^{\gamma}w=a^{\ast},$ then by squeezing $\lim_{x\longrightarrow 0}\left|x\right|^{\gamma}u=a^{\ast}$; or $\lim_{r\longrightarrow 0}r^{S_{i}}v=k_{i}^{\prime}>0,$ with $i=1$ or $2$, then also $\lim_{r\longrightarrow 0}r^{S_{i}}w=k_{i}^{\prime\prime}>0,$ implying that $u\asymp$ $\left|x\right|^{-S_{i}},$ then $\lim_{x\longrightarrow 0}\left|x\right|^{S_{i}}u=k_{i}>0$ from Proposition 5.11. For $\mu=\mu_{0}$ we still have a possibility that $\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}(\left|\ln\left|x\right|\right|^{-\frac{2}{p}})v(\left|x\right|)=\ell_{1}>0$, and $\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{-\frac{2}{p}}w(\left|x\right|)=\ell_{2}>0$, then we get (2.17) from Proposition 5.12.
(ii) Let $u$ be a solution in $\mathbb{R}^{N}\backslash\overline{B_{r_{0}}}.$ Then from Theorem 4.6, $\lim_{r\longrightarrow\infty}r^{\gamma}v=a^{\ast}=\lim_{r\longrightarrow\infty}r^{\gamma}w,$ thus $\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{\gamma}u=a^{\ast}$.
(iii) Let $u$ be a solution in $\mathbb{R}^{N}\backslash\left\{0\right\}$. Then $v\leq u\leq w\leq cv$ in $\mathbb{R}^{N}\backslash\left\{0\right\}$. From Theorem 4.6, all the behaviours of radial solutions in $\mathbb{R}^{N}\backslash\left\{0\right\}$ are still described. Either $\mu>\mu_{0}$ and $r^{\gamma}v\equiv r^{\gamma}v=a^{\ast},$ thus $\left|x\right|^{\gamma}u\equiv a^{\ast}$. Or $\mu>\mu_{0}$, and $\lim_{r\longrightarrow 0}r^{S_{1}}v=k_{i}^{\prime}>0$, and $\lim_{r\longrightarrow\infty}r^{\gamma}v=a^{\ast}$; then $\lim_{x\longrightarrow 0}\left|x\right|^{S_{1}}u=k_{i}>0$ from Proposition 5.11, and $\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{\gamma}u=a^{\ast}$. For given $k_{1}>0,$ there exists a unique function $u$ satisfying these two conditions. Indeed if $\widetilde{u}$ is another such solution, then $(1+\varepsilon)\widetilde{u}$ is a supersolution, and near 00 and $\infty$,∞ , it is greater than $u,$ then $(1+\varepsilon)\widetilde{u}\geq u;$ then $\widetilde{u}\geq u,$ and $\widetilde{u}=u.$ By uniqueness $u$ is radial. Finally if $\mu=\mu_{0},$ then $\lim_{x\rightarrow 0}\left|x\right|^{\frac{N-p}{p}}\left|\ln\left|x\right|\right|^{-\frac{2}{p}}u(x)=\ell>0$ from Proposition 5.12, and $\lim_{\left|x\right|\longrightarrow\infty}\left|x\right|^{\gamma}u(x)=a^{\ast}.$ By comparison as above, it is unique and radial.
@@ 7.1.1 Behaviour near the fixed point $\mathbf{M}_{0}$
are the roots of the trinom $T(\lambda);$ and the product of the roots is negative, thus $\mathbf{M}_{0}$ is a saddle point. The eigenvectors can be computed easily.
@@ 7.1.2 Behaviour near the fixed points $\mathbf{A}_{1}$,$\mathbf{A}_{2}$
$\det\left(\begin{array}[]{cc}pS_{i}-N+p-\lambda&-1\\ 0&(q+1-p)(\gamma-S_{i})-\lambda\end{array}\right)=0,$
@@ 7.2 Proofs and comments
then $\mathcal{T}_{1}$ lies in $\mathcal{R}_{1}$ as $t\rightarrow\infty$, then it stays in it, and necessarily converges to $\mathbf{A}_{1}$ (even in the case $\rho_{1}=\eta_{1}$). Therefore there is a unique trajectory, namely $\mathcal{T}_{1}$, joining $\mathbf{A}_{1}$ to $\mathbf{M}_{0}$, then $u$ satisfies (4.13).
$\bullet$ The trajectory $\mathcal{T}_{3}$ starts from $\mathbf{M}_{0}$ in the region $\mathcal{R}_{3};$ and $\mathcal{R}_{3}$ is positively invariant: indeed the vector field is entering $\mathcal{R}_{3}$ at any point or the curve $\mathcal{C}_{2}$, except possibly at the singular point $(0,\left|\mu\right|)$ when $\mu<0$ and $p>2$. But no trajectory with a point in $\mathcal{R}_{3}$ can pass by this point, since such trajectories satisfy $G<0$, from then $\mathcal{T}_{3}$ stays in $\overline{\mathcal{R}_{3}}$. Consider any corresponding solution $u(r)=U(t)$ defined on an maximal interval $(-\infty,T)$. Suppose that $V$ is bounded; then $F(G)$ is bounded, so $\left|G\right|$ is bounded, then $(G,V)$ converges to some $(\ell,\Lambda)$, then $(G_{t},V_{t})$ has a finite limit. It cannot happen that $(\ell,\Lambda)=(0,-\mu)$, because we have seen that the trajectories passing by this point are contained in the set $G<0$. Then $T=\infty$, thus $(G,V)$ converges to a fixed point; it is impossible, since there is no other fixed point in $\overline{\mathcal{R}_{3}}$. Then $T$ is finite, and $\lim_{t\longrightarrow T}V=\infty;$ if $\lambda=\lim_{t\longrightarrow T}G$ is finite, and $\lim_{t\longrightarrow T}\frac{V_{t}}{V}=c>0$, which contradicts $\lim_{t=T}V=\infty$. Then $\lim_{t\longrightarrow T}G=-\infty$. It implies that the function $u$, such that $u\sim_{r\longrightarrow 0}u^{\ast}$decreasing as $r\rightarrow 0$, has a minimum point, and then is increasing to $\infty$ as $r\rightarrow e^{T}$. Namely we obtain a solution $u$, satisfying (4.17), often called large solution.
@@
for some $\tau\in\left(0,1\right),$ thus $F_{1}(G_{1}+\overline{G})\sim_{t\longrightarrow-\infty}c\overline{G}^{2}$ with $c=\frac{1}{2}\frac{p}{p-1}S_{1}^{2-p},$ and
$\overline{G}_{t}=c\overline{G}^{2}(1+O(\overline{G}))-V.$
@@
In the general case $p>1$, we can solve the equation (7.13) by quadrature, and get
@@
Acknowledgment    H. Chen is supported by NSFC of China, No. 12071189, 12361043, by Jiangxi Province Science Fund No. 20232ACB201001.